# Stability of the Kim–Milman flow map

### v1

@@ Abstract
In this short note, we characterize stability of the Kim–Milman flow map—also known as the probability flow ODE—with respect to variations in the target measure. Rather than the Wasserstein distance, we show that stability holds with respect to the relative Fisher information.
@@ 1 Introduction
In general, there are infinitely many maps which transport a fixed source distribution $\rho$ to a target distribution $\mu$, both in $\mathcal{P}(\mathbb{R}^{d})$. Let $T_{\bullet}$ denote a method of generating such transport maps; thus, $T_{\bullet}^{\rho\to\mu}$ is a transport map from $\rho$ to $\mu$ for every $\mu\in\mathcal{P}(\mathbb{R}^{d})$, meaning that for $X\sim\rho$, $T_{\bullet}^{\rho\to\mu}(X)\sim\mu$. A property of fundamental interest for such a method is its stability with respect to variations in the target measure: how much do the transport maps vary if the target measures vary? That is, for another target measure $\nu$ and under a broad class of assumptions, we want to understand inequalities of the form
@@
In this note, we study stability properties of a different transport map called the Kim–Milman (reverse) heat flow map [13], which has recently gained popularity in the machine learning literature under the moniker “probability flow ODE” [12] due to its efficacy in generative modeling tasks. Unlike optimal or entropic transport maps, we stress that this transport map is defined dynamically. As a brief description, let $\gamma$ denote the $d$-dimensional standard Gaussian distribution, and let $\mu$ be another $d$-dimensional probability distribution; we want to study the stability of the following (reverse) ODE system
@@
where $\rm{FI}(\nu\,\|\,\mu)\coloneqq\|\nabla\log(\nu/\mu)\|^{2}_{L^{2}(\nu)}$, where the underlying constant is explicit. We then show how our analysis can be used to prove bounds of the form
@@
where $\operatorname{FI}_{\infty}(\nu\,\|\,\mu)\coloneqq{\rm{ess\,sup}}_{\nu}\|\nabla\log(\nu/\mu)\|^{2}$. We consider as applications the case where $\mu$ is a perturbation of a strongly log-concave measure (leveraging recent results by [3, 26]) and when $\mu$ has an asymptotically positive convexity profile (as introduced by [7]). We stress that, at present, these cases are not covered by stability results for (entropic) optimal transport maps.
@@ 2 Background
2.1 Probability flow ODE
We now review the derivation of the probability flow ODE, also known as the reverse heat flow due to Kim and Milman [13].
@@ 2.1 Probability flow ODE
which describes the evolution of marginal law of the ODE system
@@
Let $T_{\rm KM}^{\mu,T}$ be the flow map corresponding the ODE system (2.2); thus, $(T_{\rm KM}^{\mu,T})_{\sharp}q_{T}=\mu$. Finally, we let $T_{\rm KM}^{\mu}\coloneqq\lim_{T\to\infty}T_{\rm KM}^{\mu,T}$ denote the Kim–Milman map.
@@ 2.2 Related work
Most recently, this argument has been pushed by Letrouit and Mérigot in [17] to show that, so long as $\mu,\nu$ have compact support and $\rho$ has density bounded above and below on a compact convex subset of $\mathbb{R}^{d}$, then
@@
We note that this same result is true over Riemannian manifolds [14]. On the other hand, it is not possible to establish Hölder stability in general with an exponent better than $1/2$, due to the counterexample in [11].
@@
where the sole requirement is that all three measures have compact support (say in a ball of radius $R>0$). Though, the results of [9] are more general. For example, if one of the entropic transport maps is bi-Lipschitz (which can be ensured in certain situations [6]), then they are able to recover the results of [19] by taking $\varepsilon\searrow 0$.
@@ 3 Main results
We are interested in understanding the dynamics of the system (2.2). To this end, let $(X_{t})_{t\in[0,T]}$ (resp. $(Y_{t})_{t\in[0,T]}$) be the reverse dynamics from $\gamma$ to $\mu$ (resp. $\gamma$ to $\nu$) for $0\ll T<+\infty$ with $q^{\mu}_{t}\coloneqq{\rm{Law}}(X_{t})$ (resp. $q^{\nu}_{t}\coloneqq{\rm{Law}}(Y_{t})$). Thus, $T_{\rm KM}^{\mu}(x)$ is the terminal point $X_{T}$ of the ODE (2.2) with $T\to\infty$ and initialized at $x$, and similarly for $T_{\rm KM}^{\nu}(x)$. For a rigorous justification of this procedure, see, e.g., [13, 20]. Our calculations will require that for all $s\geq 0$
$\nabla^{2}\log Q_{s}\Bigl[\frac{\mu}{\gamma}\Bigr](\cdot)=I+\nabla^{2}\log q_{s}^{\mu}(\cdot)\preceq{\theta}_{s}I\,.$ ($\Theta$)
Our main results will be of the form
@@
where we place assumptions on $\mu$, and the omitted constant will depend explicitly on these assumptions. We can view this result as a strengthening of the classical transport-information inequality. For instance, an application of the HWI inequality (see [10]) for $\mu$ which is $\alpha$-strongly log-concave yields
$W_{2}^{2}(\mu,\nu)\leq\alpha^{-2}\,{\rm FI}(\nu\,\|\,\mu)\,$, (3.1)
where $W_{2}^{2}(\mu,\nu)$ denotes the squared $2$-Wasserstein distance between $\mu$ and $\nu$. However, it follows from a trivial coupling argument and Corollary 3.4 that
@@ 3.1 Main computation
In this section, we prove the following proposition.
Proposition 3.1.
Let $\mu$ be such that ($\Theta$) is satisfied for some constants $(\theta_{s})_{s\geq 0}$. Then, for any $\nu$ such that $\rm{FI}(\nu\,\|\,\mu)<\infty$ and $T\geq 0$,
$(\mathbb{E}\|X_{T}-Y_{T}\|^{2})^{1/2}$ $\leq\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,\Lambda_{T}\coloneqq\sqrt{{\rm FI}(\nu\,\|\,\mu)}\int_{0}^{T}\exp\Bigl(\int_{0}^{T-s}(3\theta_{u}-1)\,\mathrm{d}u\Bigr)\,\mathrm{d}s\,.$
Our main theorem is then the following, which follows as an immediate corollary of Proposition 3.1 by taking the appropriate limits.
@@
Suppose the conditions of Proposition 3.1 are satisfied, and write $\Lambda_{\infty}\coloneqq\lim_{T\to\infty}\Lambda_{T}$. Then,
$\|T_{\rm KM}^{\mu}-T_{\rm KM}^{\nu}\|_{L^{2}(\gamma)}\leq\Lambda_{\infty}\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,.$
To prove our main proposition, we require the following lemma due to Wibisono [26].
Lemma 3.3 ([26, Theorem 4(i)]).
@@
Proof. [Proof of Proposition 3.1] To start, we compute
@@
${\rm FI}(q_{t}^{\nu}\,\|\,q_{t}^{\mu})$ $\leq\exp\Bigl(-2\int_{0}^{t}(-2\theta_{u}+{1})\,\mathrm{d}u\Bigr)\,{\rm FI}(q_{0}^{\nu}\,\|\,q_{0}^{\mu})$
$=\exp\Bigl(-2\int_{0}^{t}(-2\theta_{u}+{1})\,\mathrm{d}u\Bigr)\,{\rm FI}(\nu\,\|\,\mu)\,.$
@@
$(\mathbb{E}\|X_{t}-Y_{t}\|^{2})^{1/2}\leq\sqrt{{\rm FI}(\nu\,\|\,\mu)}\int_{0}^{t}c_{T-s}\exp\Bigl(\int_{s}^{t}{\theta_{T-r}}\,\mathrm{d}r\Bigr)\,\mathrm{d}s\,.$
Taking $t=T$ and writing out $c_{T-s}$, our full inequality is
$(\mathbb{E}\|X_{T}-Y_{T}\|^{2})^{1/2}$ $\leq\sqrt{{\rm FI}(\nu\,\|\,\mu)}\int_{0}^{T}e^{-(T-s)}\exp\Bigl(\int_{0}^{T-s}2\theta_{u}\,\mathrm{d}u\Bigr)\exp\Bigl(\int_{s}^{T}{\theta_{T-r}}\,\mathrm{d}r\Bigr)\,\mathrm{d}s$
$=\sqrt{{\rm FI}(\nu\,\|\,\mu)}\int_{0}^{T}e^{-(T-s)}\exp\Bigl(3\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s\,$,
where the final step follows from change of variables. ∎
3.2 Corollaries
We now instantiate Proposition 3.1 with some cases of interest.
3.2.1 Strong log-concavity
As a warm-up, suppose that $\mu$ is $\alpha$-strongly log-concave, i.e., there exists $\alpha>0$ such that $0\prec\alpha I\preceq-\nabla^{2}\log\mu$. In this case, it is well-known that (see, e.g., [20])
${\theta}_{u}=\frac{1-\alpha}{\alpha\,(\exp(2u)-1)+1}\,.$
Incorporating this bound into Proposition 3.1 and carrying out the integration yields the following corollary. As this is a special case of the next section, we omit the computation.
Corollary 3.4.
Suppose that $\mu$ is $\alpha$-strongly convex. Then
$\|T_{\rm{KM}}^{\mu}-T_{\rm{KM}}^{\nu}\|_{L^{2}(\gamma)}\leq\alpha^{-1}\,\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,.$
@@ 3.2.2 Strong log-concavity with log-Lipschitz perturbations
We now suppose our main target measure is of the form $\mu\propto\exp(-V+H)$ where $V$ is $\alpha$-strongly convex and $H$ is a (smooth) $L$-Lipschitz perturbation. In this setting, a recent result of [3] showed that
@@
Carrying out the algebra, the complete stability bound in this setting is given below. Note that the special case $L=0$ recovers Corollary 3.4.
@@
Suppose $\mu\propto\exp(-V-H)$ where $V$ is $\alpha$-strongly convex and $H$ is an $L$-Lipschitz function. Then
$\|T_{\rm{KM}}^{\mu}-T_{\rm{KM}}^{\nu}\|_{L^{2}(\gamma)}\leq\frac{1}{\alpha}\exp\Bigl(\frac{3L^{2}}{2\alpha}+\frac{6L}{\sqrt{\alpha}}\Bigr)\,\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,.$
Proof.  Letting $b=\exp(2(T-s))-1$, carrying out the integration yields
$\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u=-\frac{1}{2}\log\Bigl(\frac{1+\alpha b}{1+b}\Bigr)+\frac{bL^{2}}{2\,(1+\alpha b)}+\frac{2bL}{1+\alpha b}\sqrt{\alpha+b^{-1}}\,.$
Another change of variables with $r=(b+1)^{-1/2}=\exp(-(T-s))$ yields, for the full integral,
$\int_{0}^{T}\exp(-(T-s))\exp\Bigl(3\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s$
@@
$\int_{0}^{T}\exp(-(T-s))\exp\Bigl(3\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s$
$\qquad\qquad\leq\exp\Bigl(\frac{3L^{2}}{2\alpha}+\frac{6L}{\sqrt{\alpha}}\Bigr)I(T,\alpha)$
$\qquad\qquad\coloneqq\exp\Bigl(\frac{3L^{2}}{2\alpha}+\frac{6L}{\sqrt{\alpha}}\Bigr)\int_{\exp(-T)}^{1}(r^{2}+\alpha\,(1-r^{2}))^{-3/2}\,\mathrm{d}r\,.$
Performing the integration in closed form, it is easy to see that $\lim_{T\to\infty}I(T,\alpha)=\alpha^{-1}$. This concludes the proof, as
$\int_{0}^{T}\exp(-(T-s))\exp\Bigl(3\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s$ $\leq\frac{1}{{\alpha}}\exp\Bigl(\frac{3L^{2}}{2\alpha}+\frac{6L}{\sqrt{\alpha}}\Bigr)\,.$
@@
$\|T_{\rm{KM}}^{\mu}-T_{\rm{KM}}^{\nu}\|_{L^{2}(\gamma)}$ $\leq\frac{1}{\alpha}\exp\Bigl(C\,\frac{\beta^{2}\max_{k}\|m_{k}\|^{2}}{\alpha}\Bigr)\,\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,$,
for some universal constant $C>0$.

### v2

@@ Abstract
In this short note, we characterize stability of the Kim–Milman flow map—also known as the probability flow ODE—with respect to variations in the target measure in relative Fisher information.
@@ 1 Introduction
In general, there are infinitely many maps which transport a fixed source distribution $\rho$ to a target distribution $\mu$, where both measures are in $\mathcal{P}(\mathbb{R}^{d})$. Let $T_{\bullet}$ denote a method of generating such transport maps; thus, $T_{\bullet}^{\rho\to\mu}$ is a transport map from $\rho$ to $\mu$ for every $\mu\in\mathcal{P}(\mathbb{R}^{d})$, meaning that for $X\sim\rho$, $T_{\bullet}^{\rho\to\mu}(X)\sim\mu$. A property of fundamental interest for such a method is its stability with respect to variations in the target measure: how much do the transport maps vary if the target measures vary? That is, for another target measure $\nu$ and under a broad class of assumptions, we want to understand inequalities of the form
@@
In this note, we study stability properties of a different transport map called the Kim–Milman (reverse) heat flow map [18], which has recently gained popularity in mathematics [29, 20, 28, 26, 3, 13, 7, 35, 14] as well as in the machine learning literature under the moniker “probability flow ODE” [37, 17] due to its efficacy in generative modeling tasks. Unlike optimal or entropic transport maps, we stress that this transport map is defined dynamically. As a brief description, let $\gamma$ denote the $d$-dimensional standard Gaussian distribution, and let $\mu$ be another $d$-dimensional probability distribution; we want to study the stability of the following (reverse) ODE system
@@
where $\rm{FI}(\nu\,\|\,\mu)\coloneqq\|\nabla\log(\nu/\mu)\|^{2}_{L^{2}(\nu)}$, where the underlying constant is explicit; see Theorem 3.1. We then show how our analysis can be used to prove bounds of the form
@@
where $\operatorname{FI}_{\infty}(\nu\,\|\,\mu)\coloneqq{\rm{ess\,sup}}_{\nu}\|\nabla\log(\nu/\mu)\|^{2}$; see Theorem 4.3. We consider as applications the case where $\mu$ is a perturbation of a strongly log-concave measure (leveraging recent results by [3, 38]) and when $\mu$ has an asymptotically positive convexity profile (as introduced by [9]). We stress that, at present, these cases are not covered by stability results for (entropic) optimal transport maps.
For future work, it would be natural to extend our results to measures which live on Riemannian manifolds [13] or to other flow maps (e.g., the flow map induced by general stochastic interpolants [25, 1] or the Schrödinger bridge flow map [21, 36, 33]).
Acknowledgments
We thank Zhou Fan and Daniel Lacker for helpful discussions, as well as Katharina Eichinger for pointing out a technical flaw in our original proof, and the referees for providing valuable feedback. AAP thanks the Foundations of Data Science at Yale University for support. MSZ is funded by an NSERC CGS-D award.
@@ 2 Background
2.1 Kim–Milman flow map
We now sketch the derivation of the reverse heat flow due to Kim and Milman [18], also known as the probability flow ODE.
@@ 2.1 Kim–Milman flow map
which describes the evolution of the marginal law of the ODE system
@@
Let $T_{\rm KM}^{\mu,T}$ be the flow map corresponding to the ODE system (2.2); thus, $(T_{\rm KM}^{\mu,T})_{\sharp}q_{T}=\mu$. Finally, we let $T_{\rm KM}^{\mu}\coloneqq\lim_{T\to\infty}T_{\rm KM}^{\mu,T}$ denote the Kim–Milman map.
@@ 2.2 Related work
Using different arguments, Delalande and Merigot [10] showed that, so long as $\mu,\nu$ have compact support and $\rho$ has density bounded above and below on a compact convex subset of $\mathbb{R}^{d}$, then
@@
Letrouit and Mérigot in [22] pushed these arguments to accommodate more general source measures, and [19] proposed an extension to Riemannian manifolds. On the other hand, it is not possible to establish Hölder stability in general with an exponent better than $1/2$, due to the counterexample in [16].
@@
where the sole requirement is that all three measures have compact support (say in a ball of radius $R>0$). However, the results of [11] are more general. For example, if one of the entropic transport maps is bi-Lipschitz (which can be ensured in certain situations [6]), then they are able to recover the results of [27] by taking $\varepsilon\searrow 0$.
@@ 3 Main results
We are interested in understanding the dynamics of the system (2.2). To this end, let $(X_{t})_{t\in[0,T]}$ (resp. $(Y_{t})_{t\in[0,T]}$) be the reverse dynamics from $q_{T}^{\mu}$} to $\mu$ (resp. $q_{T}^{\nu}$} to $\nu$) for $0\ll T<+\infty$ with $q^{\mu}_{{T-t}}\coloneqq{\rm{Law}}(X_{t})$ (resp. $q^{\nu}_{{T-t}}\coloneqq{\rm{Law}}(Y_{t})$). Thus, $T_{\rm KM}^{\mu}(x)$ is the terminal point $X_{T}$ of the ODE (2.2) with $T\to\infty$ and initialized at $x$, and similarly for $T_{\rm KM}^{\nu}(x)$. For a rigorous justification of this procedure, see, e.g., [18, 28]. Our calculations will require that for all $s\geq 0$
$\nabla^{2}\log Q_{s}\Bigl[\frac{\mu}{\gamma}\Bigr](\cdot)=I+\nabla^{2}\log q_{s}^{\mu}(\cdot)\preceq{\theta}_{s}I$ ($\Theta$)
for some constants $(\theta_{s})_{s\geq 0}$. Our main results will be of the form
@@
where the omitted constants will explicitly depend on assumptions on $\mu$.
@@ 3.1 Main computation
Since we only place assumptions on $\mu$, even the existence of $T_{\rm KM}^{\nu}$ is non-trivial. To avoid technical complications, we assume that $T_{\rm KM}^{\mu}$, $T_{\rm KM}^{\nu}$ are well-defined in the sense that $T_{\rm KM}^{\mu,T}\to T_{\rm KM}^{\mu}$ uniformly on compact sets and similarly for $T_{\rm KM}^{\nu}$. [^1] This is stronger than the convergence established in [28]. For more complete discussions, see [18, 28].
In this section, we prove our main theorem.
@@
Let $\mu$ be such that ($\Theta$) is satisfied for some constants $(\theta_{s})_{s\geq 0}$. Assume
$\Lambda_{\infty}\coloneqq\lim_{T\to\infty}\Lambda_{T}<\infty\,,\qquad\text{where}\penalty 10000\ \Lambda_{T}\coloneqq\int_{0}^{T}\exp\Bigl(\int_{0}^{t}(3\theta_{s}-1)\,\mathrm{d}s\Bigr)\,\mathrm{d}t\,.$
Then, for any $\nu$ such that $\rm{FI}(\nu\,\|\,\mu)<\infty$,
$\|T_{\rm KM}^{\mu}-T_{\rm KM}^{\nu}\|_{L^{2}(\gamma)}\leq\Lambda_{\infty}\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,.$ (3.1)
To prove our main theorem, we require the following lemma due to Wibisono [38].
Lemma 3.2 ([38, Theorem 4(i)]).
@@
Proof. [Proof of Theorem 3.1] To start, we compute
@@
${\rm FI}(q_{t}^{\nu}\,\|\,q_{t}^{\mu})$ $\leq\exp\Bigl(-2\int_{0}^{t}(-2\theta_{u}+{1})\,\mathrm{d}u\Bigr)\,{\rm FI}(q_{0}^{\nu}\,\|\,q_{0}^{\mu})=\exp\Bigl(-2\int_{0}^{t}(-2\theta_{u}+{1})\,\mathrm{d}u\Bigr)\,{\rm FI}(\nu\,\|\,\mu)\,.$
@@
$(\mathbb{E}\|X_{t}-Y_{t}\|^{2})^{1/2}$ $\leq{\exp\Bigl(\int_{0}^{t}\theta_{T-s}\,\mathrm{d}s\Bigr)\,(\mathbb{E}\|X_{0}-Y_{0}\|^{2})^{1/2}}$
$\qquad+\sqrt{{\rm FI}(\nu\,\|\,\mu)}\int_{0}^{t}c_{T-s}\exp\Bigl(\int_{s}^{t}{\theta_{T-r}}\,\mathrm{d}r\Bigr)\,\mathrm{d}s\,.$
Let $Z\sim\gamma$ and $X_{0}\coloneqq e^{-T}\bar{X}+\sqrt{1-e^{-2T}}\,Z$, $Y_{0}\coloneqq e^{-T}\bar{Y}+\sqrt{1-e^{-2T}}\,Z$, where $\bar{X}\sim\mu$, $\bar{Y}\sim\nu$ are coupled so that $\|\bar{X}-\bar{Y}\|_{L^{2}}=W_{2}(\mu,\nu)$. This produces a valid coupling $X_{0}\sim q_{T}^{\mu}$, $Y_{0}\sim q_{T}^{\nu}$. Moreover, $X_{0}\to Z$ almost surely as $T\to\infty$, and similarly $Y_{0}\to Z$. By the assumption that the Kim–Milman maps are well-defined, it holds that $T_{\rm KM}^{\mu,T}(X_{0})\to T_{\rm KM}^{\mu}(Z)$ and $T_{\rm KM}^{\nu,T}(Y_{0})\to T_{\rm KM}^{\nu}(Z)$. By Fatou’s lemma and the above bound,
$\|T_{\rm KM}^{\mu}(Z)-T_{\rm KM}^{\nu}(Z)\|_{L^{2}}$ $\leq\liminf_{T\to\infty}\|T_{\rm KM}^{\mu,T}(X_{0})-T_{\rm KM}^{\nu,T}(Y_{0})\|_{L^{2}}$
$\leq\liminf_{T\to\infty}\Bigl[\exp\Bigl(\int_{0}^{T}(\theta_{t}-1)\,\mathrm{d}t\Bigr)\,W_{2}(\mu,\nu)$
$\qquad{}+\sqrt{{\rm FI}(\nu\,\|\,\mu)}\int_{0}^{T}c_{T-s}\exp\Bigl(\int_{s}^{T}{\theta_{T-r}}\,\mathrm{d}r\Bigr)\,\mathrm{d}s\Bigr]\,.$
Now, we argue that the first term vanishes. We can safely assume $\Lambda_{\infty}<\infty$, or else the statement we wish to prove is trivial. Then, $\Lambda_{\infty}\geq\int_{0}^{\infty}\exp(3\int_{0}^{T}(\theta_{u}-1)\,\mathrm{d}u)\,\mathrm{d}T$ which implies that $\exp(\int_{0}^{T}(\theta_{u}-1)\,\mathrm{d}u)\to 0$ as $T\to\infty$ along a subsequence, as desired.
Now taking $t=T$ and writing out $c_{T-s}$, our proof concludes by taking a limit
${\frac{\|T_{\rm KM}^{\mu}-T_{\rm KM}^{\nu}\|_{L^{2}(\gamma)}}{\sqrt{{\rm FI}(\nu\,\|\,\mu)}}}$ $\leq{\limsup_{T\to\infty}}\int_{0}^{T}e^{-(T-s)}\exp\Bigl(\int_{0}^{T-s}2\theta_{u}\,\mathrm{d}u\Bigr)\exp\Bigl(\int_{s}^{T}{\theta_{T-r}}\,\mathrm{d}r\Bigr)\,\mathrm{d}s$
$={\limsup_{T\to\infty}}\int_{0}^{T}e^{-(T-s)}\exp\Bigl(3\int_{0}^{T-s}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s$
$={\limsup_{T\to\infty}\int_{0}^{T}\exp\Bigl(\int_{0}^{s}(3\theta_{u}-1)\,\mathrm{d}u\Bigr)\,\mathrm{d}s\,,}$
where we apply a change of variables. ∎
@@ 3.1.1 Strong log-concavity with log-Lipschitz perturbations
We first suppose our main target measure is of the form $\mu\propto\exp(-V+H)$ where $V$ is $\alpha$-strongly convex and $H$ is a (smooth) $L$-Lipschitz perturbation. In this setting, a recent result of [3] showed that
@@
Carrying out the algebra, the complete stability bound in this setting is given below.
@@
Suppose $\mu\propto\exp(-V{+}H)$ where $V$ is $\alpha$-strongly convex and $H$ is an $L$-Lipschitz function. Then (3.1) holds with
$\Lambda_{\infty}=\frac{1}{\alpha}\exp\Bigl(\frac{3L^{2}}{2\alpha}+\frac{6L}{\sqrt{\alpha}}\Bigr)\,.$
If $\mu$ is only $\alpha$-strongly log-concave (i.e., $H=0$), then $\Lambda_{\infty}=\alpha^{-1}$.
Remark 3.4.
One can view Corollary 3.3 as a strengthening of the classical transport-information inequality. For instance, an application of the HWI inequality (see [15]) for $\mu$ which is $\alpha$-strongly log-concave yields
$W_{2}^{2}(\mu,\nu)\leq\alpha^{-2}\,{\rm FI}(\nu\,\|\,\mu)\,$, (3.4)
where $W_{2}^{2}(\mu,\nu)$ denotes the squared $2$-Wasserstein distance between $\mu$ and $\nu$ [30]. However, it follows from a trivial coupling argument and Corollary 3.3 that
Proof. [Proof of Corollary 3.3] Letting $b=\exp(2{s})-1$, carrying out the integration yields
$\int_{0}^{s}\theta_{u}\,\mathrm{d}u=-\frac{1}{2}\log\Bigl(\frac{1+\alpha b}{1+b}\Bigr)+\frac{bL^{2}}{2\,(1+\alpha b)}+\frac{2bL}{1+\alpha b}\sqrt{\alpha+b^{-1}}\,.$
Another change of variables with $r=(b+1)^{-1/2}=\exp(-{s})$ yields, for the full integral,
$\int_{0}^{T}\exp(-{s})\exp\Bigl(3\int_{0}^{{s}}\theta_{u}\,\mathrm{d}u\Bigr)\,\mathrm{d}s$
@@
${\Lambda_{T}}$ $\leq\exp\Bigl(\frac{3L^{2}}{2\alpha}+\frac{6L}{\sqrt{\alpha}}\Bigr)I(T,\alpha)\coloneqq\exp\Bigl(\frac{3L^{2}}{2\alpha}+\frac{6L}{\sqrt{\alpha}}\Bigr)\int_{\exp(-T)}^{1}(r^{2}+\alpha\,(1-r^{2}))^{-3/2}\,\mathrm{d}r\,.$
Performing the integration in closed form, it is easy to see that $\lim_{T\to\infty}I(T,\alpha)=\alpha^{-1}$. This concludes the proof. ∎
@@
$\|T_{\rm{KM}}^{\mu}-T_{\rm{KM}}^{\nu}\|_{L^{2}(\gamma)}$ $\leq\frac{1}{\alpha}\exp\Bigl(O\Bigl(\frac{\beta^{2}\max_{k}\|m_{k}\|^{2}}{\alpha}\Bigr)\Bigr)\,\sqrt{{\rm FI}(\nu\,\|\,\mu)}\,.$